A 68 kg bungee jumper is standing on a platform 46
Answer #1

h0 = 46m

Ha = 9m

Hb = 31m

k = 66 N/m^2 and g =9.80m/s^2 , m= 68 kg

a) Speed after falling 9m

1/2 m(Va^2) + mg(Ha) = mg(h0)

Va = Root (2g (h0-Ha))

= Root [2*9.8m/s^2 (46m-9m)]

Va = 26.9 m/s

b) Speed after falling 31m

mg(h0) =mg(Hb)+k[(Ha)-(Hb)]^2/2 +mVb^2/2

68kg*9.8m/s^2*46m = 68kg * 9.8m/s^2 (31m) + 66N/m[9-31]^2 /2 + 68*Vb^2 /2

30654.4 = 9996+31944+ 34Vb^2

Vb^2 = (30654.4-9996-31944)/34 = -331.92

Vb = 18.21 m/s

c) sketch are attached below